# Joint moments of characteristic polynomials from the orthogonal and unitary symplectic groups

### v1

@@ 2 Convergence of joint moments of power sums
Abusing notation we will always write $\mathbb{P}$ for the underlying laws of these random variables, even if not necessarily defined on the same probability space, and $\mathbb{E}$ for the corresponding expectations. Finaly define, for $k\in\mathbb{N}$, the negative power sums:
@@
for some $r\geq 1$, and $h_{1},\ldots,h_{m}\in[0,\infty)$ are such that $\sum_{k=1}^{m}kh_{k}<r$, then we also have convergence of joint moments:
@@
so that $M_{N}^{(k)}=\mathsf{Z}_{N,l}^{(k)}+\mathsf{Y}_{N,l}^{(k)}$ and $M^{(k)}=\mathsf{Z}_{l}^{(k)}+\mathsf{Y}_{l}^{(k)}$ for each $l,N\in\mathbb{N}$. Throughout this section we also denote $Q=\sum_{k}kh_{k}$ whenever the choice of $h_{1},\ldots,h_{m}\in[0,\infty)$ is clear. We begin with the following lemma.
@@
$\limsup_{N\to\infty}\mathbb{E}\left[\mathsf{Y}_{l,N}^{(1)}\right]=\limsup_{N\to\infty}\bigg{\{}\mathbb{E}\left[M_{N}^{(1)}\right]-\mathbb{E}\left[\mathsf{Z}_{l,N}^{(1)}\right]\bigg{\}}\leq\mathbb{E}\left[M^{(1)}\right]-\mathbb{E}\left[\mathsf{Z}_{l}^{(1)}\right]=\mathbb{E}\left[\mathsf{Y}_{l}^{(1)}\right].$
@@
$\int_{0}^{\infty}\mathbb{P}(\mathsf{T}_{l}\geq t)\mathrm{d}t=\mathbb{E}[\mathsf{T}_{l}]\leq\liminf_{j\to\infty}\mathbb{E}\left[\mathsf{Y}_{l,N_{j}}^{(1)}\right]\leq\mathbb{E}[\mathsf{Y}_{l}^{(1)}]=\int_{0}^{\infty}\mathbb{P}(\mathsf{Y}_{l}^{(1)}>t)\mathrm{d}t,$
@@
for almost every $t\in(0,\infty)$. But then the function $t\mapsto\mathbb{P}(\mathsf{T}_{l}\geq t)$ can have countably many discontinuities, so that $\mathbb{P}(\mathsf{T}_{l}>t)=\mathbb{P}(\mathsf{Y}_{l}^{(1)}>t)$ for almost every $t\in(0,\infty)$, hence showing that $\mathsf{T}_{l}$ must have the same law as $\mathsf{Y}_{l}^{(1)}$. In particular, we conclude that every subsequence of $\{\mathsf{Y}_{l,N}^{(1)}\}_{N\geq 1}$ has a further subsequence that converges in law to $\mathsf{Y}_{l}^{(1)}$, which then implies the desired convergence for the full sequence. This completes the proof of the lemma. ∎
@@
where $S\stackrel{{\scriptstyle\mathrm{def}}}{{=}}\sum_{k=1}^{m}h_{k}$, and the sum is over all subsets of $\{1,\ldots,m\}$ including $\emptyset$.
@@
Define, for each $k\in\{1,\ldots,m\}$, $p_{k}=\frac{h_{k}}{S}$. It is easy to see that
@@
The result now follows by taking products over $k=1,\dots,m$, expanding the product, and using the triangle inequality for $\mathrm{L}^{S}\big{(}\mu_{N}^{(a,b)}\big{)}$. ∎
@@
$T_{N,J}^{(l)}=\mathbb{E}\left[\prod_{k\in J}\left(\mathsf{Z}_{N,l}^{(k)}\right)^{h_{k}}\prod_{k\notin J}\left(\mathsf{Y}_{N,l}^{(k)}\right)^{h_{k}}\right].$
@@
Combining this with the uniform integrability given by applying concavity of $t\mapsto t^{1/k}$ to each multiplicand, and making use of the assumption (15), we get the convergence of averages. Moving on, note once again utilizing positivity, we have
$\liminf_{N\to\infty}\mathbb{E}\left[\prod_{j=1}^{m}(M_{N}^{(k)})^{h_{k}}\right]\geq\liminf_{N\to\infty}\mathbb{E}\left[\prod_{k=1}^{m}\left(\sum_{i=1}^{l}\frac{1}{\lambda_{i,N}^{k}}\right)^{h_{k}}\right]=\mathbb{E}\left[\prod_{k=1}^{m}\left(\sum_{i=1}^{l}\frac{1}{\gamma_{i}^{k}}\right)^{h_{k}}\right].$ (21)
@@
$\liminf_{N\to\infty}\mathbb{E}\left[\prod_{k=1}^{m}(M_{N}^{(k)})^{h_{k}}\right]\geq\mathbb{E}\left[\prod_{j=1}^{m}(M^{(k)})^{h_{k}}\right].$
@@
$\limsup_{l\to\infty}\limsup_{N\to\infty}T_{N,J}^{(l)}=0.$
@@
$T_{N,J}^{(l)}\leq\prod_{k\in J}\mathbb{E}\bigg{[}(\mathsf{Z}_{N,l}^{(k)})^{Q/k}\bigg{]}^{\frac{kh_{k}}{Q}}\prod_{k\notin J}\mathbb{E}\bigg{[}(\mathsf{Y}_{N,l}^{(k)})^{Q/k}\bigg{]}^{\frac{kh_{k}}{Q}}.$
The claim now follows by observing that for each $k$, we have $(\mathsf{Z}_{N,l}^{(k)})^{Q/k}\leq(\mathsf{Z}_{N,l}^{(1)})^{Q}$ and $(\mathsf{Y}_{N,l}^{(k)})^{Q/k}\leq(\mathsf{Y}_{N,l}^{(1)})^{Q}$, and also that by means of Lemma 2.2 along with assumption (15),
@@
Hence, combining this with (20) and (21) we get the convergence of joint moments. For the joint convergence in law, pick, under the given assumptions, $h_{1},\ldots,h_{m}\in\mathbb{R}_{+}$ such that $Q=\sum_{k}kh_{k}<r$. First, note that by a combination of Markov’s inequality and Prokhorov’s theorem, the random vectors
@@
$\int_{\mathbb{R}_{+}^{m}}\mathbb{P}(Y_{1}\geq t_{1},\ldots,$ $Y_{m}\geq t_{m})\mathrm{d}\mathbf{t}=\mathbb{E}\left[Y_{1}\cdots Y_{m}\right]=\lim_{N\to\infty}\mathbb{E}\left[(M_{N}^{(1)})^{h_{1}}\cdots(M_{N}^{(m)})^{h_{m}}\right]$
@@ 2.1 Application to Jacobi ensemble
$\mathcal{Z}_{N}^{a,b}=\frac{1}{N!}\prod_{j=1}^{N}\frac{\Gamma(a+j)\Gamma(b+j)\Gamma(1+j)}{\Gamma(a+b+N+j)\Gamma(2)}.$
For $k\in\mathbb{N}$, define the negative power sums
$\mathsf{M}_{N}^{(k)}\stackrel{{\scriptstyle\mathrm{def}}}{{=}}\sum_{j=1}^{N}\frac{1}{\big{(}\lambda_{j,N}\big{)}^{k}},$
where $\lambda_{j,N}=N^{2}\mathsf{x}_{j,N}$, and $\big{(}\mathsf{x}_{1,N},\ldots,\mathsf{x}_{N,N}\big{)}$ are distributed according to $\mu_{N}^{(a,b)}$. The $N^{2}$-scaling here is the so-called hard-edge scaling for the Jacobi ensemble. Note that $\mathsf{M}_{N}^{(k)}$ depends on $a$ and $b$ but we suppress it from the notation. We also define their Bessel analogues:
$\mathsf{M}^{(k)}\stackrel{{\scriptstyle\mathrm{def}}}{{=}}\sum_{j=1}^{\infty}\frac{1}{\big{(}\mathfrak{B}_{a}^{j}\big{)}^{k}}.$
Note that $\mathsf{M}^{(k)}$ depends on $a$ (there is no parameter $b$ here) but we suppress it from the notation. We then have the following result, where here, and throughout the rest of the paper, for any positive real number $x$, we define
@@
$\big{(}\mathsf{M}_{N}^{(1)},\ldots,\mathsf{M}_{N}^{(m)}\big{)}\stackrel{{\scriptstyle\mathrm{d}}}{{\longrightarrow}}\big{(}\mathsf{M}^{(1)},\ldots,\mathsf{M}^{(m)}\big{)},$
@@
$\mathbb{E}_{N}^{(a,b)}\left[\prod_{k=1}^{m}(\mathsf{M}_{N}^{(k)})^{h_{k}}\right]\longrightarrow\mathbb{E}^{(a)}\left[\prod_{k=1}^{m}(\mathsf{M}^{(k)})^{h_{k}}\right]<\infty,\ \ \mathrm{as}\ N\to\infty.$
@@
$\lim_{N\to\infty}\mathbb{E}_{N}^{(a,b)}\Big{[}\mathsf{M}_{N}^{(1)}\Big{]}=\frac{1}{a}=\mathbb{E}\big{[}\mathsf{M}^{(1)}\big{]},$
where the first equality follows from Aomoto’s integral formula (see, for instance, [32]), and the second equality is seen by first writing the average as the one-dimensional integral $\int_{0}^{\infty}x^{-1}\mathsf{K}_{a}^{\mathrm{Bes}}(x,x)\mathrm{d}x$, thanks to the determinantal property of the process $\{\mathfrak{B}_{a}^{k}\}_{k=1}^{\infty}$, and computing this integral explicitly using formulae for Bessel functions (a roundabout way to obtain the desired formula would be to combine results of [9, 5]). Finally, as a consequence of [35, Proposition 4.1], we get that whenever $\alpha\in\mathbb{N}$ with $\alpha<a+1$, we have
$\sup_{N}\mathbb{E}_{N}^{(a,b)}\Big{[}\big{(}\mathsf{M}_{N}^{(1)}\big{)}^{\alpha}\Big{]}<\infty,$
@@
Using standard arguments, see, for instance, [34, Proposition 5.1.2], this function can be written as:
@@
and the $\big{\{}\mathcal{P}_{j}^{(a,b)}\big{\}}_{j=1}^{\infty}$ are polynomials indexed by their degree defined by the orthonormality condition:
@@
$\frac{C(\alpha)}{N^{\kappa}}\prod_{j\in A}\int\frac{1}{y^{\sum_{J\in A_{j}}m_{J}}}K_{N}^{(a,b)}(y,y)\mathrm{d}y\leq\frac{C(\alpha)}{N^{\kappa}}\prod_{j\in A}N^{2\sum_{j\in J}m_{J}}=C(\alpha)N^{2\alpha l-\kappa}$
@@
that holds for all $1\leq\theta<a+1$ (see the proof of Proposition 2.2 in [40]). Combining these observations, one then has
$\mathbb{E}\left[\mathrm{e}_{l}\bigg{(}\frac{1}{\mathsf{x}_{1}^{(N)}},\ldots,\frac{1}{\mathsf{x}_{N}^{(N)}}\bigg{)}^{\alpha}\right]\leq C(\alpha)\sum_{\sum_{J\in[N]^{(l)}}m_{J}=\alpha}N^{2\alpha l-\kappa((m_{J})_{J_{\in}[N]^{(l)}})}.$
@@
$C(\alpha,l)\,\binom{N}{\kappa}$
@@
Let $a>0$, b>1b>-1, and $m\in\mathbb{N}$. Then there exist random vectors $(\mathfrak{J}_{N,1}^{(a,b)},\ldots,\mathfrak{J}_{N,l}^{(a,b)})_{N\geq 1}$ and $(\mathfrak{J}_{1}(a),\ldots,\mathfrak{J}_{l}(a))$ defined on a common probability space such that
$(\mathfrak{e}_{N,1}^{(a,b)},\ldots,\mathfrak{e}_{N,k}^{(a,b)})\stackrel{{\scriptstyle\mathrm{d}}}{{=}}(\mathfrak{J}_{N,1}^{(a,b)},\ldots,\mathfrak{J}_{N,k}^{(a,b)})$
@@
$(\mathfrak{e}_{1}(a),\ldots,\mathfrak{e}_{k}(a))\stackrel{{\scriptstyle\mathrm{d}}}{{=}}(\mathfrak{J}_{1}(a),\ldots,\mathfrak{J}_{k}(a)).$
@@
$(N^{-2}\mathfrak{J}_{N,1},\ldots,N^{-2k}\mathfrak{J}_{N,k})\longrightarrow(\mathfrak{J}_{1},\ldots,\mathfrak{J}_{k}),\ \ \textnormal{as }N\to\infty.$
@@
Define $\mathbb{H}_{+}(\infty)=\lim_{\leftarrow}\mathbb{H}_{+}(N)$, the projective limit of the $\mathbb{H}_{+}(N)$ under these maps, and
@@
Let $\mu\in\mathcal{M}_{\mathrm{p}}^{\mathrm{Perm}}$ and $\mathbf{H}$ having law $\mu$, such that for some $m\in\mathbb{N}$, $\det(\mathbf{H}_{k})\in\mathrm{L}^{p}(\mu)$ for every $k\in[m]$, where $\mathbf{H}_{k}=\mathfrak{P}_{k}^{\infty}(\mathbf{H})$. Then, if we denote by $(\mathsf{x}_{1}^{(N)},\ldots,\mathsf{x}_{N}^{(N)})$ the eigenvalues of $\mathbf{H}_{N}$, we have
@@
as $N\to\infty$, $\mu$-almost surely and in $\mathrm{L}^{p}(\mu)$, for some random variables $\{\mathsf{a}_{k}\}_{k=1}^{m}$. Furthermore, for $h_{1},\ldots,h_{k}\in\mathbb{N}$, we have the formula, with $Q=\sum_{k}kh_{k}$,
@@
$\left(N^{-1}\sum_{i=1}^{N}\frac{1}{\mathsf{y}_{i}^{(N)}},N^{-2}\sum_{i=1}^{N}\frac{1}{(\mathsf{y}_{i}^{(N)})^{2}},\dots,N^{-m}\sum_{i=1}^{N}\frac{1}{(\mathsf{y}_{i}^{(N)})^{m}}\right)\stackrel{{\scriptstyle\mathrm{d}}}{{\longrightarrow}}\left(\sum_{i=1}^{\infty}\frac{1}{\mathfrak{B}_{a}^{i}},\sum_{i=1}^{\infty}\frac{1}{(\mathfrak{B}_{a}^{i})^{2}},\dots,\sum_{i=1}^{\infty}\frac{1}{(\mathfrak{B}_{a}^{i})^{k}}\right)$

### v2

@@ 2 Convergence of joint moments of power sums
Abusing notation we will always write $\mathbb{P}$ for the underlying laws of these random variables, even if not necessarily defined on the same probability space, and $\mathbb{E}$ for the corresponding expectations. Finally define, for $k\in\mathbb{N}$, the negative power sums:
@@
for some $r\geq 1$, and $h_{1},\ldots,h_{m}\in[0,\infty)$ are such that $\sum_{k=1}^{m}kh_{k}<r$, then we also have convergence of joint moments as $N\rightarrow\infty$:
@@
so that $M_{N}^{(k)}=\mathsf{Z}_{N,l}^{(k)}+\mathsf{Y}_{N,l}^{(k)}$ and $M^{(k)}=\mathsf{Z}_{l}^{(k)}+\mathsf{Y}_{l}^{(k)}$ for each $l,N\in\mathbb{N}$. We begin with the following lemma.
@@
$\limsup_{N\to\infty}\mathbb{E}\left[\mathsf{Y}_{N,l}^{(1)}\right]=\limsup_{N\to\infty}\bigg\{\mathbb{E}\left[M_{N}^{(1)}\right]-\mathbb{E}\left[\mathsf{Z}_{N,l}^{(1)}\right]\bigg\}\leq\mathbb{E}\left[M^{(1)}\right]-\mathbb{E}\left[\mathsf{Z}_{l}^{(1)}\right]=\mathbb{E}\left[\mathsf{Y}_{l}^{(1)}\right].$
@@
$\int_{0}^{\infty}\mathbb{P}(\mathsf{T}_{l}\geq t)\mathrm{d}t=\mathbb{E}[\mathsf{T}_{l}]\leq\liminf_{j\to\infty}\mathbb{E}\left[\mathsf{Y}_{N_{j},l}^{(1)}\right]\leq\mathbb{E}[\mathsf{Y}_{l}^{(1)}]=\int_{0}^{\infty}\mathbb{P}(\mathsf{Y}_{l}^{(1)}>t)\mathrm{d}t,$
@@
for almost every $t\in(0,\infty)$. But then the function $t\mapsto\mathbb{P}(\mathsf{T}_{l}\geq t)$ can have countably many discontinuities, so that $\mathbb{P}(\mathsf{T}_{l}>t)=\mathbb{P}(\mathsf{Y}_{l}^{(1)}>t)$ for almost every $t\in(0,\infty)$, hence showing that $\mathsf{T}_{l}$ must have the same law as $\mathsf{Y}_{l}^{(1)}$. In particular, we conclude that every subsequence of $\{\mathsf{Y}_{N,l}^{(1)}\}_{N\geq 1}$ has a further subsequence that converges in law to $\mathsf{Y}_{l}^{(1)}$, which then implies the desired convergence for the full sequence. This completes the proof of the lemma. ∎
@@
where $S\stackrel{{\scriptstyle\mathrm{def}}}{{=}}\max(\sum_{k=1}^{m}h_{k},1)$ with $h_{1},\ldots,h_{m}\in[0,\infty)$, and the sum is over all subsets of $\{1,\ldots,m\}$ including $\emptyset$.
@@
First, consider the case $\sum_{k=1}^{m}h_{k}\geq 1$. Define, for each $k\in\{1,\ldots,m\}$, $p_{k}=\frac{h_{k}}{S}$. It is easy to see that
@@
The inequality now follows by taking products over $k=1,\dots,m$, expanding the product, and using the triangle inequality for $\mathrm{L}^{S}$. If, on the other hand, $\sum_{k=1}^{m}h_{k}<1$, then $h_{j}\in[0,1)$ for $j=1,\ldots,m$, so that the inequality follows immediately by using the sub-additivity of the map $t\mapsto t^{h_{k}}$. ∎
@@
$\mathsf{T}_{N,J}^{(l)}=\mathbb{E}\left[\prod_{k\in J}\left(\mathsf{Z}_{N,l}^{(k)}\right)^{h_{k}}\prod_{k\notin J}\left(\mathsf{Y}_{N,l}^{(k)}\right)^{h_{k}}\right].$
@@
Combining this with the uniform integrability given by applying the subadditivity of $t\mapsto t^{1/k}$ on $0,\infty)$ to each multiplicand, and making use of the assumption (15), we obtain the convergence of averages. Moving on, note once again utilizing positivity, we have
$\liminf_{N\to\infty}\mathbb{E}\left[\prod_{j=1}^{m}(M_{N}^{(k)})^{h_{k}}\right]\geq\liminf_{N\to\infty}\mathbb{E}\left[\prod_{k=1}^{m}\left(\sum_{i=1}^{l}\frac{1}{\lambda_{i,N}^{k}}\right)^{h_{k}}\right]=\mathbb{E}\left[\prod_{k=1}^{m}\left(\sum_{i=1}^{l}\frac{1}{\gamma_{i}^{k}}\right)^{h_{k}}\right].$
@@
$\liminf_{N\to\infty}\mathbb{E}\left[\prod_{k=1}^{m}(M_{N}^{(k)})^{h_{k}}\right]\geq\mathbb{E}\left[\prod_{k=1}^{m}(M^{(k)})^{h_{k}}\right].$ (21)
@@
$\limsup_{l\to\infty}\limsup_{N\to\infty}\mathsf{T}_{N,J}^{(l)}=0.$ (22)
@@
$\mathsf{T}_{N,J}^{(l)}\leq\prod_{k\in J}\mathbb{E}\bigg[(\mathsf{Z}_{N,l}^{(k)})^{Q/k}\bigg]^{\frac{kh_{k}}{Q}}\prod_{k\notin J}\mathbb{E}\bigg[(\mathsf{Y}_{N,l}^{(k)})^{Q/k}\bigg]^{\frac{kh_{k}}{Q}}.$
The claim (2.3) now follows by observing that for each $k$, we have $(\mathsf{Z}_{N,l}^{(k)})^{Q/k}\leq(\mathsf{Z}_{N,l}^{(1)})^{Q}$ and $(\mathsf{Y}_{N,l}^{(k)})^{Q/k}\leq(\mathsf{Y}_{N,l}^{(1)})^{Q}$, and also that by means of Lemma 2.2 along with assumption (15),
@@
Then, combining with (20), (22) and Lemma 2.3,
$\limsup_{N\to\infty}\mathbb{E}\left[\prod_{j=1}^{m}(M_{N}^{(k)})^{h_{k}}\right]\leq\mathbb{E}\left[\prod_{k=1}^{m}\left(M^{(k)}\right)^{h_{k}}\right].$ (23)
By (21) and (23), we have the convergence of joint moments.
For the joint convergence in law, pick, under the given assumptions, $h_{1},\ldots,h_{m}\in\mathbb{R}_{+}$ such that $Q=\sum_{k}kh_{k}<r$. First, note that by a combination of Markov’s inequality and Prokhorov’s theorem, the random vectors
@@
$\int_{\mathbb{R}_{+}^{m}}\mathbb{P}(Y_{1}\geq t_{1},\ldots,$ $Y_{m}\geq t_{m})\mathrm{d}\mathbf{t}=\mathbb{E}\left[Y_{1}\cdots Y_{m}\right]=\lim_{j\to\infty}\mathbb{E}\left[(M_{N_{j}}^{(1)})^{h_{1}}\cdots(M_{N_{j}}^{(m)})^{h_{m}}\right]$
@@ 2.1 Application to Jacobi ensemble
$\mathcal{Z}_{N}^{a,b}=\frac{1}{N!}\prod_{j=1}^{N}\frac{\Gamma(a+j)\Gamma(b+j)\Gamma(1+j)}{\Gamma(a+b+N+j)}.$
For $k\in\mathbb{N}$,recall that we defined earlier the negative power sums
where $\lambda_{j,N}=N^{2}\mathsf{x}_{j}^{(N)}$, and $\big(\mathsf{x}_{1}^{(N)},\ldots,\mathsf{x}_{N}^{(N)}\big)$ are distributed according to $\mu_{N}^{(a,b)}$. The $N^{2}$-scaling here is the so-called hard-edge scaling for the Jacobi ensemble. In what follows, we suppress the parameters $a,b$ from the notation whenever their values are clear. Also recall that we defined their Bessel analogues:
$\mathfrak{p}_{m}(a)\stackrel{{\scriptstyle\mathrm{def}}}{{=}}\sum_{j=1}^{\infty}\frac{1}{\big(\mathfrak{B}_{a}^{j}\big)^{m}}.$
We then have the following result, where here, and throughout the rest of the paper, for any positive real number $x$, we define
@@
$\big(\mathfrak{p}_{1,N}^{(a,b)},\ldots,\mathfrak{p}_{m,N}^{(a,b)}\big)\stackrel{{\scriptstyle\mathrm{d}}}{{\longrightarrow}}\big(\mathfrak{p}_{1}(a),\ldots,\mathfrak{p}_{m}(a)\big),$
@@
$\mathbb{E}_{N}^{(a,b)}\left[\prod_{k=1}^{m}(\mathfrak{p}_{k,N})^{h_{k}}\right]\longrightarrow\mathbb{E}\left[\prod_{k=1}^{m}(\mathfrak{p}_{k}(a))^{h_{k}}\right]<\infty,\ \ \mathrm{as}\ N\to\infty.$
@@
$\lim_{N\to\infty}\mathbb{E}_{N}^{(a,b)}\Big[\mathfrak{p}_{1,N}\Big]=\frac{1}{a}=\mathbb{E}\big[\mathfrak{p}_{1}(a)\big],$
where the first equality follows from Aomoto’s integral formula (see, for instance, [31]), and the second equality is seen by first writing the average as the one-dimensional integral $\int_{0}^{\infty}x^{-1}\mathsf{K}_{a}^{\mathrm{Bes}}(x,x)\mathrm{d}x$, thanks to the determinantal property of the process $\{\mathfrak{B}_{a}^{k}\}_{k=1}^{\infty}$, and computing this integral explicitly using formulae for Bessel functions (a roundabout way to obtain the desired formula would be to combine results of [9, 5]). Finally, as a consequence of [34, Proposition 4.1], we obtain that whenever $\alpha\in\mathbb{N}$ with $\alpha<a+1$, we have
$\sup_{N}\mathbb{E}_{N}^{(a,b)}\Big[\big(\mathfrak{p}_{1,N}\big)^{\alpha}\Big]<\infty,$
@@
Using standard arguments, see, for instance, [33, Proposition 5.1.2], this function can be written as:
@@
and the $\big\{\mathcal{P}_{j}^{(a,b)}\big\}_{j=1}^{\infty}$ are polynomials indexed by their degrees defined by the orthonormality condition:
@@
$\frac{C(\alpha)}{\binom{N}{\kappa}}\prod_{j\in A}\int\frac{1}{y^{\sum_{J\in A_{j}}m_{J}}}K_{N}^{(a,b)}(y,y)\mathrm{d}y\leq\frac{\tilde{C}(a,b,\alpha)}{\binom{N}{\kappa}}\prod_{j\in A}N^{2\sum_{j\in J}m_{J}}=\frac{\tilde{C}(a,b,\alpha)}{\binom{N}{\kappa}}N^{2\alpha l}$
@@
that holds for all $1\leq\theta<a+1$ (see the proof of Proposition 2.2 in [39]), and $\tilde{C}(a,b,\alpha)=C(\alpha)\sup_{1\leq\theta<a+1}C(a,b,\theta)$. Combining these observations, one then has
$\mathbb{E}_{N}^{(a,b)}\left[\mathrm{e}_{l}\bigg(\frac{1}{\mathsf{x}_{1}^{(N)}},\ldots,\frac{1}{\mathsf{x}_{N}^{(N)}}\bigg)^{\alpha}\right]\leq\tilde{C}(a,b,\alpha)\sum_{\sum_{J\in[N]^{(l)}}m_{J}=\alpha}N^{2\alpha l}\binom{N}{\kappa((m_{J})_{J_{\in}[N]^{(l)}})}^{-1}.$
@@
$\tilde{C}(\alpha,l)\,\binom{N}{\kappa}$
@@
Let $a>0$, b>1b>-1, and $m\in\mathbb{N}$. Then there exist random vectors $(\mathfrak{J}_{N,1}^{(a,b)},\ldots,\mathfrak{J}_{N,m}^{(a,b)})_{N\geq 1}$ and $(\mathfrak{J}_{1}(a),\ldots,\mathfrak{J}_{m}(a))$ defined on a common probability space such that
$(\mathfrak{e}_{N,1}^{(a,b)},\ldots,\mathfrak{e}_{N,m}^{(a,b)})\stackrel{{\scriptstyle\mathrm{d}}}{{=}}(\mathfrak{J}_{N,1}^{(a,b)},\ldots,\mathfrak{J}_{N,m}^{(a,b)})$
@@
$(\mathfrak{e}_{1}(a),\ldots,\mathfrak{e}_{m}(a))\stackrel{{\scriptstyle\mathrm{d}}}{{=}}(\mathfrak{J}_{1}(a),\ldots,\mathfrak{J}_{m}(a)).$
@@
$(N^{-2}\mathfrak{J}_{N,1},\ldots,N^{-2m}\mathfrak{J}_{N,m})\longrightarrow(\mathfrak{J}_{1},\ldots,\mathfrak{J}_{m}),\ \ \textnormal{as }N\to\infty.$
@@
[…]
@@
Let $\mu\in\mathcal{M}_{\mathrm{p}}^{\mathrm{Perm}}$ and $\mathbf{H}$ having law $\mu$, such that for some $m\in\mathbb{N}$, $p\in[1,\infty)$ such that $\det(\mathbf{H}_{k})\in\mathrm{L}^{p}(\mu)$ for every $k\in[m]$, where $\mathbf{H}_{k}=\mathfrak{P}_{k}^{\infty}(\mathbf{H})$. Then, if we denote by $(\mathsf{x}_{1}^{(N)},\ldots,\mathsf{x}_{N}^{(N)})$ the eigenvalues of $\mathbf{H}_{N}$, we have
@@
as $N\to\infty$, $\mu$-almost surely and in $\mathrm{L}^{p}(\mu)$, for some random variables $\{\mathsf{a}_{k}\}_{k=1}^{m}$. Furthermore, for $h_{1},\ldots,h_{m}\in\mathbb{N}\cup\{0\}$ with $h_{m}\neq 0$, we have the formula, with $Q=\sum_{k}kh_{k}$,
@@
$\left(N^{-1}\sum_{i=1}^{N}\frac{1}{\mathsf{y}_{i}^{(N)}},N^{-2}\sum_{i=1}^{N}\frac{1}{(\mathsf{y}_{i}^{(N)})^{2}},\dots,N^{-m}\sum_{i=1}^{N}\frac{1}{(\mathsf{y}_{i}^{(N)})^{m}}\right)\stackrel{{\scriptstyle\mathrm{d}}}{{\longrightarrow}}\left(\sum_{i=1}^{\infty}\frac{1}{\mathfrak{B}_{a}^{i}},\sum_{i=1}^{\infty}\frac{1}{(\mathfrak{B}_{a}^{i})^{2}},\dots,\sum_{i=1}^{\infty}\frac{1}{(\mathfrak{B}_{a}^{i})^{m}}\right)$